OpenStudy (mantar0078):

Please help Simplify. Identify any x-values for which the expression is undefined. 3x+4 over 3x^2+x−4 Answers: the expression is undefined for x=-4/3 and x=1 1 over x−1 the expression is undefined for x=3/4 and x=-1/2 1 over x+3 the expression is undefined for x=3/4 and x=-1/2 3 over x+1 the expression is undefined for x=-4/3 and x=1 x−1

Directrix (directrix):

First, factor the expression 3x^2+x−4 so that we can find values of x that make the denominator equal to zero.

Directrix (directrix):

3x^2+x−4 = ( x - 1) * ( ? + ?)

Directrix (directrix):

@mantar0078 ^^^^ Need the factors.

OpenStudy (mantar0078):

How do I find the factors?

Directrix (directrix):

Do you recall the factoring of polynomials from Algebra One? Does this look familiar? Fill in for the second factor. 3x^2+x−4 = ( x - 1) * ( ? + ?)